How is the answer to part b: 8/15? I am calculating 7/15...

An urn contains 2 blue marbles and 4 red marbles. One marble is​ removed, its color​ noted, and not replaced. A second marble is removed and its color is noted.

​(a) What is the probability that both marbles are blue​? red​?

​(b) What is the probability that exactly one marble is blue​?
Answer #1

Here 2 blue and 4 red marbles so we are to find

P( exactly one marble is blue)=

(2/6) *(4/5) +(4/6) *(2/5) =16/30=8/15